29. Which of the ordered pairs is a solution for the equation 2x – 5y = 10?
A) (0, 2)
B) (–5, 0)
C) (–5, –2)
D) (0, –2)
Thanks for your help!!!

Answer by checkley71(8403)

2X-5Y=10
2*0-5*2=10
0-10=10
-10=10 NOT A SOLUTION
2*-5-5*0=10
-10-0=10
-10=10 NOT A SOLUTION
2*-5-5*-2=10
-10+10=10
0=10 NOT A SOLUTION
2*0-5*-2=10
0+10=10
10=10 THIS IS A SOLUTION D)(0,-2)